# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v3

@@ Abstract
Zekun Zhou1, Xiaocheng Feng1,2thanks: Corresponding Author, Lei Huang1, Xiachong Feng3, Ziyun Song1, Ruihan Chen1, Liang Zhao1, Weitao Ma1, Yuxuan Gu1, Baoxin Wang4, Dayong Wu4, Guoping Hu4, Ting Liu1, Bing Qin1,2 1 Harbin Institute of Technology, Harbin, China  2 Peng Cheng Laboratory, Shenzhen, China 3 The University of Hong Kong, China  4 iFLYTEK Research, China {zkzhou,xcfeng,lhuang,zysong,rhchen,lzhao,wtma,yxgu,tliu,qinb}@ir.hit.edu.cn fengxc@hku.hk{bxwang2,dywu2,gphu}@iflytek.com
@@ 1 Introduction
Research is creative and systematic work aimed at expanding knowledge and driving civilization’s development (Eurostat, 2018). Researchers typically identify a topic, review relevant literature, synthesize existing knowledge, and formulate hypothesis, which are validated through theoretical and experimental methods. Findings are then documented in manuscripts that undergo peer review before publication (Benos et al., 2007; Boyko et al., 2023). However, this process is resource-intensive, requiring specialized expertise and posing entry barriers for researchers (Blaxter et al., 2010).
In recent years, artificial intelligence (AI) technologies, represented by large language models (LLMs), have experienced rapid development  (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Yang et al., 2024a; DeepSeek-AI et al., 2024; Guo et al., 2025). These models exhibit exceptional capabilities in text understanding, reasoning, and generation (Schaeffer et al., 2023). In this context, AI is increasingly involving the entire research pipeline (Messeri and Crockett, 2024), sparking extensive discussion about its implications for research  (Hutson, 2022; Williams et al., 2023; Morris, 2023; Fecher et al., 2023). Moreover, following the release of ChatGPT, approximately 20% of academic papers and peer-reviewed texts in certain fields have been modified by LLMs (Liang et al., 2024a, b). A study also reveals that 81% of researchers integrate LLMs into their workflows (Liao et al., 2024).
@@
Comparing with Existing Surveys Although Luo et al. (2025) reviews the application of AI in research, it predominantly focuses on LLMs, while neglecting the knowledge synthesis that precedes hypothesis generation and the theoretical validation of hypothesis. Other surveys concentrate on more specific areas, such as paper recommendation (Beel et al., 2016; Bai et al., 2019; Kreutz and Schenkel, 2022), scientific literature review (Altmami and Menai, 2022), hypothesis generation (Kulkarni et al., 2025), scientific claim verification (Vladika and Matthes, 2023; Dmonte et al., 2024), theorem proving (Li et al., 2024e), manuscript writing (Li and Ouyang, 2024), and peer review  (Lin et al., 2023a; Kousha and Thelwall, 2024). Additionally, certain surveys emphasize the application of AI in scientific domains (Zheng et al., 2023b; Zhang et al., 2024d; Gridach et al., 2025).
@@
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8em,font=,, where level=2text width=13em,font=,, where level=3text width=11.5em,font=,, where level=4text width=13em,font=,, [ A survey of AI for Research, ver [ Hypothesis
Formulation  (§2) [ Knowledge Synthesis  (§2.1) [ Research Paper
Recommendation  (§2.1.1) [ E.g.,  ComLittee Kang et al. (2023) , PaperWeaver Lee et al. (2024) ,leaf, text width=34.5em ] ] [ Systematic
Literature Review  (§2.1.2) [ E.g.,  AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) ,leaf, text width=34.5em ] ] ] [ Hypotheses Generation  (§2.2) [ E.g.,  SciMON  Wang et al. (2024c) , MOOSE  Yang et al. (2024b) , Dolphin  Yuan et al. (2025) ,leaf, text width=47.6em ] ] ] [ Hypothesis
Validation  (§3) [ Scientific
Claim Verification  (§3.1) [ E.g.,  FactKG  Kim et al. (2023) , SFAVEL Bazaga et al. (2024) , ClaimVer Dammu et al. (2024) ,leaf, text width=47.6em ] ] [ Theorem Proving  (§3.2) [ E.g.,  Dt-solver  Wang et al. (2023b) , LEGO-Prover  Wang et al. (2024b) , DeepSeek-Prover  Xin et al. (2024) ,leaf, text width=47.6em ] ] [ Experiment Validation  (§3.3) [ E.g.,  SANDBOX  Liu et al. (2023b) , CRISPR-GPT  Huang et al. (2024a) , MLR-Copilot  Li et al. (2024d) ,leaf, text width=47.6em ] ] ] [ Manuscript
Publication  (§4) [ Manuscript Writing  (§4.1) [ E.g.,  Scilit  Gu and Hahnloser (2023) , UR3WG  Shi et al. (2023) , STEP-BACK  Tang et al. (2024a) ,leaf, text width=47.6em ] ] [ Peer Review  (§4.2) [ E.g.,  SWIF2T  Chamoun et al. (2024) , GLIMPSE  Darrin et al. (2024) , MetaWriter  Sun et al. (2024) ,leaf, text width=47.6em ] ] ] ]
@@
Survey Organization We first elaborate hypothesis formulation (§2), followed by hypothesis validation (§3) and manuscript publication (§4). Additionally, we present benchmarks (§5), and tools (§6) that utilized in research. Finally, we outline challenges as well as future directions (§7). In the Appendix, we provide further discussion on open questions (§A), challenges faced in different domains (§B), discusssion about relevant ethical considerations (§C), and a comparison of capabilities among different methods (§D).
@@ 2.1 Knowledge Synthesis
Knowledge synthesis constitutes the foundational step in the research process. During this phase, researchers are required to identify and critically evaluate existing literature to establish a thorough understanding of the field. This step is pivotal for uncovering new research directions, refining methodologies, and supporting evidence-based decision-making (Asai et al., 2024). In this section, the process of knowledge synthesis is divided into two modules: Research Paper Recommendation and Systematic Literature Review.
@@ 2.1.1 Research Paper Recommendation
Research paper recommendation (RPR) identifies and recommends novel and seminal articles aligned with researchers’ interests. Prior studies have shown that recommendation systems outperform keyword-based search engines in terms of efficiency and reliability when extracting valuable insights from large-scale datasets (Bai et al., 2019). Existing methodologies are broadly categorized into four paradigms: content-based filtering, collaborative filtering, graph-based approaches, and hybrid systems (Beel et al., 2016; Li and Zou, 2019; Bai et al., 2019; Shahid et al., 2020). Recent advancements propose multi-dimensional classification frameworks based on data source utilization  (Kreutz and Schenkel, 2022).
Recent trends in research suggest a decline in publication volumes related to RPR  (Sharma et al., 2023), alongside an increasing focus on user-centric optimizations. Existing studies emphasize the limitations of traditional paper-centric interaction models and advocate for more effective utilization of author relationship graphs  (Kang et al., 2023). Multi-stage recommendation architectures that integrate diverse methodologies have been shown to achieve superior performance  (Pinedo et al., 2024; Stergiopoulos et al., 2024). Visualization techniques that link recommended papers to users’ publication histories via knowledge graphs  (Kang et al., 2022) and LLMs-powered comparative analysis frameworks  (Lee et al., 2024) represent emerging directions for enhancing interpretability and contextual relevance.
@@ 2.1.2 Systematic Literature Review
Systematic literature review (SLR) constitutes a rigorous and structured methodology for evaluating and integrating prior research on a specific topic  (Webster and Watson, 2002; Zhu et al., 2023; Bolaños et al., 2024). In contrast to single-document summaries (Elhadad et al., 2005), SLR entails synthesizing information across multiple related scientific documents (Altmami and Menai, 2022). SLR can further be divided into two stages: outline generation and full-text generation (Shao et al., 2024; Agarwal et al., 2024b; Block and Kuckertz, 2024).
Outline generation, especially structured outline generation, is highlighted by recent studies as a pivotal factor in enhancing the quality of surveys.  Zhu et al. (2023) demonstrated that hierarchical frameworks substantially enhance survey coherence. AutoSurvey  (Wang et al., 2024e) extended conventional outline generation by recommending both sub-chapter titles and detailed content descriptions, ensuring comprehensive topic coverage. Additionally, multi-level topic generation via clustering methodologies has been proposed as an effective strategy for organizing survey structures  (Katz et al., 2024). Advanced systems such as STORM  (Shao et al., 2024) employed LLM-driven outline drafting combined with multi-agent discussion cycles to iteratively refine the generated outlines. Tree-based hierarchical architectures have gained increasing adoption in this domain. For instance, CHIME  (Hsu et al., 2024) optimized LLM-generated hierarchies through human-AI collaboration, while HiReview  (Hu et al., 2024b) demonstrated the efficacy of multi-layer tree representations for systematic knowledge organization.
Full-text generation follows the outline generation stage. AutoSurvey and  Lai et al. (2024) utilized LLMs with carefully designed prompts to construct comprehensive literature reviews step-by-step. PaperQA2  (Skarlinski et al., 2024) introduced an iterative agent-based approach for generating reviews, while STORM employed multi-agent conversation data for this purpose. LitLLM  (Agarwal et al., 2024a) and  Agarwal et al. (2024b) adopted a plan-based search enhancement strategy. KGSum  (Wang et al., 2022a) integrated knowledge graph information into paper encoding and used a two-stage decoder for summary generation. Bio-SIEVE  (Robinson et al., 2023) and  Susnjak et al. (2024) fine-tuned LLMs for automatic review generation. OpenScholar  (Asai et al., 2024) developed a pipeline that trained a new model without relying on a dedicated survey-generation model.
@@ 2.2 Hypothesis Generation
Hypothesis generation, known as idea generation, refers to the process of coming up with new concepts, solutions, or approaches. It is the most important step in driving the progress of the entire research  (Qi et al., 2023).
@@
Early work focused more on predicting relationships between concepts, because researchers believed that new concepts come from links with old concepts (Henry and McInnes, 2017; Krenn et al., 2022). As language models became more powerful (Zhao et al., 2023a), researchers are beginning to focus on open-ended idea generation  (Girotra et al., 2023; Si et al., 2024; Kumar et al., 2024). Recent advancements in AI-driven hypothesis generation highlight diverse approaches to research conceptualization. For instance, MOOSE-Chem  (Yang et al., 2024c) and IdeaSynth  (Pu et al., 2024) used LLMs to bridge inspiration-to-hypothesis transformation via interactive frameworks. The remaining research primarily falls into two areas: enhancing input data quality and improving the quality of generated hypothesis.
@@
Input data quality improvement is demonstrated by Majumder et al. (2024a); Liu et al. (2024a), who showed that LLMs can generate comprehensive hypothesis from existing academic data. Literature organization strategies have evolved through various methodologies, including triplet representations  (Wang et al., 2024c), chain-based architectures  (Li et al., 2024a), and complex database systems  (Wang et al., 2024d). Knowledge graphs emerge as critical infrastructure (Hogan et al., 2021), enabling semantic relationship mapping via subgraph identification (Buehler, 2024; Ghafarollahi and Buehler, 2024). Notably, SciMuse  (Gu and Krenn, 2024) pioneered researcher-specific hypothesis generation by constructing personalized knowledge graphs.
@@
Hypothesis quality improvement has been addressed through feedback and iteration (Rabby et al., 2025), as proposed by HypoGeniC  (Zhou et al., 2024) and MOOSE  (Yang et al., 2024b). Feedback mechanisms include direct responses to hypothesis  (Baek et al., 2024), experimental outcome evaluations  (Ma et al., 2024; Yuan et al., 2025), comparison with the existing literature (Schmidgall and Moor, 2025), and automated peer review comments  (Lu et al., 2024). FunSearch (Romera-Paredes et al., 2024) generates solutions by iteratively combining the innovative capabilities of LLM with the verification capabilities of an evaluator. Beyond iterative feedback, collaborative efforts among researchers have also been recognized, leading to multi-agent hypothesis generation approaches  (Nigam et al., 2024; Ghafarollahi and Buehler, 2024). VIRSCI  (Su et al., 2024) further optimized this process by customizing knowledge for each agent. Additionally, the Nova framework  (Hu et al., 2024a) was introduced to refine hypothesis by leveraging outputs from other research as input.
@@
Knowledge synthesis and hypothesis generation comprise the hypothesis formulation phase. Research paper recommendation supports knowledge acquisition, while systematic literature review aids organization within knowledge synthesis. Recent advances integrate LLMs  (de la Torre-López et al., 2023) to enhance knowledge integration  (Huang and Tan, 2023; Gupta et al., 2023; Kacena et al., 2024; Tang et al., 2024b). By developing a deep understanding of a domain through knowledge synthesis, researchers can identify research directions and use hypothesis generation techniques to formulate hypothesis. Additionally, the distinction between scientific discovery and hypothesis generation is discussed in  §A.

### v4

@@ Abstract
Zekun Zhou1, Xiaocheng Feng1,2thanks: Co-Corresponding Author, Lei Huang1, Xiachong Feng311footnotemark: 1, Ziyun Song1, Ruihan Chen1, Liang Zhao1, Weitao Ma1, Yuxuan Gu1, Baoxin Wang4, Dayong Wu4, Guoping Hu4, Ting Liu1, Bing Qin1,2 1 Harbin Institute of Technology  2 Peng Cheng Laboratory 3 The University of Hong Kong  4 iFLYTEK Research {zkzhou,xcfeng,lhuang,yxgu,qinb}@ir.hit.edu.cn fengxc@hku.hk{bxwang2,dywu2,gphu}@iflytek.com
@@ 1 Introduction
Research is creative and systematic work aimed at expanding knowledge and driving civilization’s development (eurostat2018measurement). Researchers typically identify a topic, review relevant literature, synthesize existing knowledge, and formulate hypothesis, which are validated through theoretical and experimental methods. Findings are then documented in manuscripts that undergo peer review before publication (benos2007ups; 2311JamesBoykoAn). However, this process is resource-intensive, requiring specialized expertise and posing entry barriers for researchers (10BlaxterHow).
In recent years, artificial intelligence (AI) technologies, represented by large language models (LLMs), have experienced rapid development  (20BrownLanguage; 2303OpenAIGPT4; 2407AbhimanyuLlama3; 2412AnYangQwen2.5; 2412DeepSeekAIDeepSeekV3; 2501deepseekAIDeepseek-r1). These models exhibit exceptional capabilities in text understanding, reasoning, and generation (23RylanSchaefferAlice). In this context, AI is increasingly involving the entire research pipeline (2024MesseriArtificial), sparking extensive discussion about its implications for research  (22HutsonCould; 2303NigelLWilliamsAlgorithmic; 2304MeredithScientists; 2306BenediktFecherFriend). Moreover, following the release of ChatGPT, approximately 20% of academic papers and peer-reviewed texts in certain fields have been modified by LLMs (2403WeixinLiangMonitoring; 2404WeixinLiangMapping). A study also reveals that 81% of researchers integrate LLMs into their workflows (2411ZhehuiLiaoLLMs).
@@
Comparing with Existing Surveys Although 25LuoZimingLLM4SR reviews the application of AI in research, it predominantly focuses on LLMs, while neglecting the knowledge synthesis that precedes hypothesis generation and the theoretical validation of hypothesis. Other surveys concentrate on more specific areas, such as paper recommendation (16beelpaper; 19XiaomeiBaiScientific; 2201ChristinKatharinaKreutzScientific), scientific literature review (22NoufIbrahimAltmamiAutomatic), hypothesis generation (kulkarni2025scientific), scientific claim verification (2305JurajVladikaScientific; 2408AlphaeusDmonteClaim), theorem proving (2404ZhaoyuLiA), manuscript writing (2404XiangciLiRelated), and peer review  (2111JialiangLinAutomated; 23KayvanKoushaArtificial). Additionally, certain surveys emphasize the application of AI in scientific domains (2310YizhenZhengLarge; 2406YuZhangA; gridach2025agentic).
@@
[…]
A survey of AI for Research
Hypothesis Formulation  (§2) Knowledge Synthesis  (§2.1) Research Paper Recommendation  (§2.1.1) E.g.,  ComLittee 2302HyeonsuB.KangComLittee , PaperWeaver 2403YoonjooLeePaperWeaver Systematic Literature Review  (§2.1.2) E.g.,  AutoSurvey  2406YidongWangAutoSurvey , OpenScholar  2411AkariAsaiOpenScholar Hypotheses Generation  (§2.2) E.g.,  SciMON  2305QingyunWangSciMON , MOOSE  2309ZonglinYangLarge , Dolphin  2501YuanJiakangDolphin Hypothesis Validation  (§3) Scientific Claim Verification  (§3.1) E.g.,  FactKG  2305JihoKimFactKG , SFAVEL 2309AdrinBazagaUnsupervised , ClaimVer 2403PreetamPrabhuSrikarDammuClaimVer Theorem Proving  (§3.2) E.g.,  Dt-solver  23DT-SolverHaimingWang , LEGO-Prover  2310HaimingWangLEGO-Prover , DeepSeek-Prover  2410HuajianXinDeepSeek-Prover Experiment Validation  (§3.3) E.g.,  SANDBOX  2305RuiboLiuTraining , CRISPR-GPT  2404KaixuanHuangCRISPRGPT , MLR-Copilot  2408RuochenLiMLRCopilot Manuscript Publication  (§4) Manuscript Writing  (§4.1) E.g.,  Scilit  2306NianlongGuSciLit , UR3WG  23ZhengliangShiTowards , STEP-BACK  2406XiangruTangStep Peer Review  (§4.2) E.g.,  SWIF2T  2405EricChamounAutomated , GLIMPSE  2406MaximeDarrinGLIMPSE , MetaWriter  24LuSunMetaWriter
@@
Survey Organization We first elaborate hypothesis formulation (§2), followed by hypothesis validation (§3) and manuscript publication (§4). Additionally, we present benchmarks (§5), and tools (§6) that utilized in research. Finally, we outline challenges as well as future directions (§7), and discusssion about relevant ethical considerations (§8). In the Appendix, we provide further discussion on open questions (§A), challenges faced in different domains (§B), and a comparison of capabilities among different methods (§C).
@@ 2.1 Knowledge Synthesis
Knowledge synthesis constitutes the foundational step in the research process. During this phase, researchers are required to identify and critically evaluate existing literature to establish a thorough understanding of the field. This step is pivotal for uncovering new research directions, refining methodologies, and supporting evidence-based decision-making (2411AkariAsaiOpenScholar). In this section, the process of knowledge synthesis is divided into two modules: Research Paper Recommendation and Systematic Literature Review.
@@ 2.1.1 Research Paper Recommendation
Research paper recommendation (RPR) identifies and recommends novel and seminal articles aligned with researchers’ interests. Prior studies have shown that recommendation systems outperform keyword-based search engines in terms of efficiency and reliability when extracting valuable insights from large-scale datasets (19XiaomeiBaiScientific). Existing methodologies are broadly categorized into four paradigms: content-based filtering, collaborative filtering, graph-based approaches, and hybrid systems (16beelpaper; 1901ZhiLiA; 19XiaomeiBaiScientific; 20shahidinsights). Recent advancements propose multi-dimensional classification frameworks based on data source utilization  (2201ChristinKatharinaKreutzScientific).
Recent trends in research suggest a decline in publication volumes related to RPR  (23RituSharmaAn), alongside an increasing focus on user-centric optimizations. Existing studies emphasize the limitations of traditional paper-centric interaction models and advocate for more effective utilization of author relationship graphs  (2302HyeonsuB.KangComLittee). Multi-stage recommendation architectures that integrate diverse methodologies have been shown to achieve superior performance  (2410IratxePinedoArZiGo; 24VaiosStergiopoulosAn). Visualization techniques that link recommended papers to users’ publication histories via knowledge graphs  (22HyeonsuBKangFrom) and LLMs-powered comparative analysis frameworks  (2403YoonjooLeePaperWeaver) represent emerging directions for enhancing interpretability and contextual relevance.
@@ 2.1.2 Systematic Literature Review
Systematic literature review (SLR) constitutes a rigorous and structured methodology for evaluating and integrating prior research on a specific topic  (02JaneWebsterAnalyzing; 2304KunZhuHierarchical; 2402FranciscoBolaArtificial). In contrast to single-document summaries (elhadad2005customization), SLR entails synthesizing information across multiple related scientific documents (22NoufIbrahimAltmamiAutomatic). SLR can further be divided into two stages: outline generation and full-text generation (2402YijiaShaoAssisting; 2412agarwalllms; 24BlockWhat).
Outline generation, especially structured outline generation, is highlighted by recent studies as a pivotal factor in enhancing the quality of surveys.  2304KunZhuHierarchical demonstrated that hierarchical frameworks substantially enhance survey coherence. AutoSurvey  (2406YidongWangAutoSurvey) extended conventional outline generation by recommending both sub-chapter titles and detailed content descriptions, ensuring comprehensive topic coverage. Additionally, multi-level topic generation via clustering methodologies has been proposed as an effective strategy for organizing survey structures  (2408UriKatzKnowledge). Advanced systems such as STORM  (2402YijiaShaoAssisting) employed LLM-driven outline drafting combined with multi-agent discussion cycles to iteratively refine the generated outlines. Tree-based hierarchical architectures have gained increasing adoption in this domain. For instance, CHIME  (2407ChaoChunHsuCHIME) optimized LLM-generated hierarchies through human-AI collaboration, while HiReview  (2410HuYuntongHireview) demonstrated the efficacy of multi-layer tree representations for systematic knowledge organization.
Full-text generation follows the outline generation stage. AutoSurvey and  2408YuxuanLaiInstruct utilized LLMs with carefully designed prompts to construct comprehensive literature reviews step-by-step. PaperQA2  (2409MichaelDSkarlinskiLanguage) introduced an iterative agent-based approach for generating reviews, while STORM employed multi-agent conversation data for this purpose. LitLLM  (2402ShubhamAgarwalLitLLM) and  2412agarwalllms adopted a plan-based search enhancement strategy. KGSum  (2209PanchengWangMulti-Document) integrated knowledge graph information into paper encoding and used a two-stage decoder for summary generation. Bio-SIEVE  (2308AmbroseRobinsonBio-SIEVE) and  2404Teosusnjakautomating fine-tuned LLMs for automatic review generation. OpenScholar  (2411AkariAsaiOpenScholar) developed a pipeline that trained a new model without relying on a dedicated survey-generation model.
@@ 2.2 Hypothesis Generation
Hypothesis generation, known as idea generation, refers to the process of coming up with new concepts, solutions, or approaches. It is the most important step in driving the progress of the entire research  (2311BiQingQiLarge).
@@
Early work focused more on predicting relationships between concepts, because researchers believed that new concepts come from links with old concepts (1708SamHenryLiterature; 2210MarioKrennPredicting). As language models became more powerful (LLMsurveyzhaoxin), researchers are beginning to focus on open-ended idea generation  (23GirotraIdeas; 2409ChengleiSiCan; 2409SandeepKumarCan). Recent advancements in AI-driven hypothesis generation highlight diverse approaches to research conceptualization. For instance, MOOSE-Chem  (2410ZonglinYangMOOSE) and IdeaSynth  (2410KevinPuIdeaSynth) used LLMs to bridge inspiration-to-hypothesis transformation via interactive frameworks. The remaining research primarily falls into two areas: enhancing input data quality and improving the quality of generated hypothesis.
@@
Input data quality improvement is demonstrated by 2402BodhisattwaPrasadMajumder; 2410HaokunLiuLiterature, who showed that LLMs can generate comprehensive hypothesis from existing academic data. Literature organization strategies have evolved through various methodologies, including triplet representations  (2305QingyunWangSciMON), chain-based architectures  (2410LongLiChain), and complex database systems  (2410WenxiaoWangSciPIP). Knowledge graphs emerge as critical infrastructure (21HoganKnowledge), enabling semantic relationship mapping via subgraph identification (2403MarkusJBuehlerAccelerating; 2409AlirezaGhafarollahiSciAgents). Notably, SciMuse  (2405XuemeiGuGeneration) pioneered researcher-specific hypothesis generation by constructing personalized knowledge graphs.
@@
Hypothesis quality improvement has been addressed through feedback and iteration (rabby2025iterative), as proposed by HypoGeniC  (2404YangqiaoyuZhouHypothesis) and MOOSE  (2309ZonglinYangLarge). Feedback mechanisms include direct responses to hypothesis  (2404JinheonBaekResearchAgent), experimental outcome evaluations  (2405PingchuanMaLLM; 2501YuanJiakangDolphin), comparison with the existing literature (schmidgall2025agentrxiv), and automated peer review comments  (2408ChrisLuThe). FunSearch (romera2024mathematical) generates solutions by iteratively combining the innovative capabilities of LLM with the verification capabilities of an evaluator. Beyond iterative feedback, collaborative efforts among researchers have also been recognized, leading to multi-agent hypothesis generation approaches  (2403HarshitNigamAcceleron; 2409AlirezaGhafarollahiSciAgents). VIRSCI  (2410HaoyangSuTwo) further optimized this process by customizing knowledge for each agent. Additionally, the Nova framework  (2410XiangHuNova) was introduced to refine hypothesis by leveraging outputs from other research as input.
@@
Knowledge synthesis and hypothesis generation comprise the hypothesis formulation phase. Research paper recommendation supports knowledge acquisition, while systematic literature review aids organization within knowledge synthesis. Recent advances integrate LLMs  (2401JosedelaTorreLopezArtificial) to enhance knowledge integration  (23HuangJingshanThe; 23GuptaRohunUtilization; 24KacenaMelissaAThe; 2412TangXuemeiAre). By developing a deep understanding of a domain through knowledge synthesis, researchers can identify research directions and use hypothesis generation techniques to formulate hypothesis. Additionally, the distinction between scientific discovery and hypothesis generation is discussed in  §A.